HHHHEEELLLPPPP.....one more please??? solve using substitution method......5x + 7y = -23 and -8x + y = 49 type an ordered pair
Answer by scott8148(6628)

y = 49 + 8x

substituting ___ 5x + 7(49 + 8x) = -23 ___ 5x + 343 + 56x = -23

solve for x, and then substitute back to find y
(use any equation that has both x and y)